# BiasLab: Toward Explainable Political Bias Detection with Dual-Axis Annotations and Rationale Indicators

### v1

@@ VII Discussion and Limitations
Despite these constraints, BiasLab offers a replicable, perception-grounded platform for evaluating media bias, explainability, and alignment between human and machine.
VIII Conclusion
BiasLab introduces a new resource for explainable political bias detection, pairing dual-axis sentiment labels with rationale indicators that surface why an article reads as left- or right-leaning. Our human-vs-LLM comparisons expose where both groups falter—especially on subtle framings—and our two baseline tasks (perception drift and rationale classification) underline the real-world difficulty of building transparent bias models. Moving forward, we plan to integrate richer contextual signals and interactive, human-in-the-loop workflows to strengthen explanation quality. By releasing the full dataset, annotation schema, and code, we invite the community to explore new methods for turning subjective bias judgments into robust, actionable insights.
Acknowledgments
This research builds upon an initial annotation methodology developed under the supervision of Dr. Dan Goldwasser during the author’s time at Purdue University. Dr. Goldwasser provided guidance and funding for the original MTurk pilot phase. The data collection, cleaning, schema formalization, full dataset release, LLM simulations, and all new analyses presented here were conducted solely by the author. We also thank Prahlad Anand for his help during the annotation phase.

### v2

@@ 9 Limitations
Acknowledgment
This research builds upon an initial annotation methodology developed under the supervision of Dr. Dan Goldwasser during the author’s time at Purdue University. Dr. Goldwasser provided guidance and funding for the original MTurk pilot phase. The data collection, cleaning, schema formalization, full dataset release, LLM simulations, and all new analyses presented here were conducted solely by the author.
Ethics Statement
This research used publicly available political news articles and involved crowd-sourced annotations via Amazon Mechanical Turk. Workers were restricted to U.S.-based accounts with a strong approval history. A qualification test was used to filter for quality. No personal or sensitive data was collected.
Appendix
Figure 7: Event distribution across issues for the 414 events covering the 900 articles
Figure 8: Hierarchical organization of bias rationale indicators used in annotation. The seven individual types are grouped into broader conceptual themes to support interpretation and modeling.
Figure 9: Batch-wise accuracy (%) of GPT-4o annotations across 30 batches (300 articles).
Table 6: Full per-class Logistic Regression results for rationale type classification. Class 1 represents positive class (presence of rationale); Class 0 is the negative class.
Rationale Type Class Precision Recall F1 Score
directional class_0 0.69 0.94 0.79
directional class_1 0.55 0.15 0.24
structural class_0 0.67 0.89 0.76
structural class_1 0.56 0.23 0.32
neutral_other class_0 0.58 0.92 0.71
neutral_other class_1 0.81 0.36 0.5